A point charge 50C is located in the XY plane at the point of position vector 0 v0= 2i + 3j . What is the electric field at the point of position vector r = 8i -5j

Grade:12

2 Answers

Deep
6 years ago
Q=50uC=50×10^-6K= 9×10^9r-r°=8i-5j-(2i+3j) =6i-8j=√6^2 +8^2 =√100 =10So R =10We know that E=kq\(R)^2 so by putting values we getE=9×10^9×50×10^-6\10^2 =4500 V\M
ankit singh
3 years ago
1 and F2 have the same magnitude and their ycomponents cancel. FQ=2. F1 . Cosθ𝚤̂=2 ×kq1Qr2.3r𝚤̂= 6kq1Qr3𝚤̂=6 ×9 ×109×5 ×2 ×10−12125𝚤̂=4.3 ×10−3𝚤̂ ( N)=4.3 𝚤